\begin{thebibliography}{10}

\bibitem{AHV95}
S.~Abiteboul, R.~Hull, and V.~Vianu.
\newblock {\em Foundations of Databases}.
\newblock Addison-Wesley, 1995.

\bibitem{ABC00}
M.~Arenas, L.~E. Bertossi, and J.~Chomicki.
\newblock Querying database repairs using logic programs with exceptions.
\newblock In {\em FQAS 2000}, pages 27--41. Physica-Verlag Heidelberg New York,
  A Springer-Verlag Company, 2000.

\bibitem{Benedikt02}
M.~Benedikt, G.~Bruns, J.~Gibson, R.~Kuss, and A.~Ng.
\newblock {Automated Update Management for XML Integrity Constraints}.
\newblock In {\em Informal Proceedings of PLAN-X Workshop}, 2002.

\bibitem{DBLP:conf/pods/Bonatti04}
P.~A. Bonatti.
\newblock On the decidability of containment of recursive datalog queries -
  preliminary report.
\newblock In A.~Deutsch, editor, {\em Proceedings of the Twenty-third ACM
  SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems, June 14-16,
  2004, Paris, France}, pages 297--306. ACM, 2004.

\bibitem{bryDeckerManthey88}
F.~Bry, H.~Decker, and R.~Manthey.
\newblock A uniform approach to constraint satisfaction and constraint
  satisfiability in deductive databases.
\newblock In {\em EDBT}, pages 488--505, 1998.

\bibitem{CCM:EROW2007}
A.~Cal{\`\i}, D.~Calvanese, and D.~Martinenghi.
\newblock Optimization of query plans in the presence of access limitations.
\newblock In {\em Proc.\ of the ICDT~2007 Workshop on Emerging Research
  Opportunities in Web Data Management (EROW~2007)}, CEUR Electronic Workshop
  Proceedings, {\upshape\protect\url{http://ceur-ws.org/}}, pages 33--47, 2007.

\bibitem{CCM:JUCS2009}
A.~Cal{\`{i}}, D.~Calvanese, and D.~Martinenghi.
\newblock {Dynamic Query Optimization under Access Limitations and
  Dependencies}.
\newblock {\em Journal of Universal Computer Science}, 15(21):33--62, 2009.

\bibitem{CM:ER2008}
A.~Cal{\`{i}} and D.~Martinenghi.
\newblock {Conjunctive Query Containment under Access Limitations}.
\newblock In {\em Proceedings of Conceptual Modeling - ER 2008, 27th
  International Conference on Conceptual Modeling, Barcelona, Spain, October
  20-24, 2008}, pages 326--340, 2008.

\bibitem{CM:ICDE2008}
A.~Cal{\`{i}} and D.~Martinenghi.
\newblock {Querying Data under Access Limitations}.
\newblock In {\em Proceedings of the 24th International Conference on Data
  Engineering, ICDE 2008, April 7-12, 2008, Canc{\'u}n, M{\'e}xico}, pages
  50--59, 2008.

\bibitem{CM:APWEB2010}
A.~Cal\`{\i} and D.~Martinenghi.
\newblock {Optimizing Query Processing for the Hidden Web (Tutorial)}.
\newblock In {\em Advances in Web Technologies and Applications, Proceedings of
  the 12th Asia-Pacific Web Conference, APWeb 2010, Busan, Korea, 6-8 April
  2010}, page 397, 2010.

\bibitem{CM:TPLP2010}
A.~Cal{\`{i}} and D.~Martinenghi.
\newblock Querying incomplete data over extended er schemata.
\newblock {\em Theory and Practice of Logic Programming}, 10(3):291--329, 2010.

\bibitem{CM:EDBT2010}
A.~Cal\`{\i} and D.~Martinenghi.
\newblock Querying the deep web (tutorial).
\newblock In {\em EDBT 2010, 13th International Conference on Extending
  Database Technology, Lausanne, Switzerland, March 22-26, 2010, Proceedings},
  pages 724--727, 2010.

\bibitem{CMC:SEBD2007}
A.~Cal{\`{i}}, D.~Martinenghi, and D.~Carbotta.
\newblock {Query optimisation for web data sources: minimisation of the number
  of accesses (Extended Abstract)}.
\newblock In {\em Proceedings of the Fifteenth Italian Symposium on Advanced
  Database Systems, SEBD 2007, 17-20 June 2007, Torre Canne, Fasano, BR,
  Italy}, pages 316--323, 2007.

\bibitem{Calvanese:1998:DQC}
D.~Calvanese, G.~{De Giacomo}, and M.~Lenzerini.
\newblock On the decidability of query containment under constraints.
\newblock In {\em Proceedings of the seventeenth ACM SIGACT-SIGMOD-SIGART
  symposium on Principles of database systems}, pages 149--158. ACM Press,
  1998.

\bibitem{CDV03}
D.~Calvanese, G.~{De Giacomo}, and M.~Vardi.
\newblock Decidable containment of recursive queries.
\newblock In D.~Calvanese, M.~Lenzerini, and R.~Motwani, editors, {\em Database
  Theory - ICDT 2003, 9th International Conference, Siena, Italy, January 8-10,
  2003, Proceedings}, volume 2572 of {\em Lecture Notes in Computer Science},
  pages 330--345. Springer, 2003.

\bibitem{Carrico95}
S.~Carrico, B.~Ewbank, T.~G. Griffin, J.~Meale, and H.~Trickey.
\newblock A tool for developing safe and efficient database transactions.
\newblock In {\em XV International Switching Symposium of the World
  Telecomminications Congress. April, 1995}, pages 173--177, 1995.

\bibitem{CCFMT:TODS2013}
I.~Catallo, E.~Ciceri, P.~Fraternali, D.~Martinenghi, and M.~Tagliasacchi.
\newblock Top-k diversity queries over bounded regions.
\newblock {\em ACM Transactions on Database Systems}, 38(2):10, 2013.
\newblock Extended version of~\cite{FMT:SIGMOD2012}.

\bibitem{DBLP:journals/tods/CeriFPT94}
S.~Ceri, P.~Fraternali, S.~Paraboschi, and L.~Tanca.
\newblock Automatic generation of production rules for integrity maintenance.
\newblock {\em ACM Transactions on Database Systems (TODS)}, 19(3):367--422,
  1994.

\bibitem{CMT:TECHREP2009}
S.~Ceri, D.~Martinenghi, and M.~Tagliasacchi.
\newblock Cost-aware rank-join algorithms.
\newblock Technical report, Politecnico di Milano, 2009.

\bibitem{DBLP:conf/vldb/CeriW90}
S.~Ceri and J.~Widom.
\newblock Deriving production rules for constraint maintainance.
\newblock In D.~McLeod, R.~Sacks-Davis, and H.-J. Schek, editors, {\em 16th
  International Conference on Very Large Data Bases, August 13-16, 1990,
  Brisbane, Queensland, Australia, Proceedings}, pages 566--577. Morgan
  Kaufmann, 1990.

\bibitem{CM:AAI2000}
H.~Christiansen and D.~Martinenghi.
\newblock {Symbolic Constraints for Meta-Logic Programming}.
\newblock {\em Applied Artificial Intelligence}, 14(4):345--367, 2000.

\bibitem{CM:LOPSTR2003}
H.~Christiansen and D.~Martinenghi.
\newblock Simplification of database integrity constraints revisited: A
  transformational approach.
\newblock In {\em Logic Based Program Synthesis and Transformation, 13th
  International Symposium LOPSTR 2003, Uppsala, Sweden, August 25-27, 2003,
  Revised Selected Papers}, volume 3018 of {\em Lecture Notes in Computer
  Science}, pages 178--197. Springer, 2004.

\bibitem{CM:FoIKS2004}
H.~Christiansen and D.~Martinenghi.
\newblock Simplification of integrity constraints for data integration.
\newblock In {\em Foundations of Information and Knowledge Systems, Third
  International Symposium, FoIKS 2004, Wilhelminenburg Castle, Austria,
  February 17-20, 2004, Proceedings}, volume 2942 of {\em Lecture Notes in
  Computer Science}, pages 31--48. Springer, 2004.

\bibitem{CM:LPAR2005}
H.~Christiansen and D.~Martinenghi.
\newblock Incremental integrity checking: Limitations and possibilities.
\newblock In {\em Logic for Programming, Artificial Intelligence, and
  Reasoning, 12th International Conference, LPAR 2005, Montego Bay, Jamaica,
  December 2-6, 2005, Proceedings}, volume 3835 of {\em Lecture Notes in
  Computer Science}, pages 712–--727. Springer, 2005.

\bibitem{CM:FI2006}
H.~Christiansen and D.~Martinenghi.
\newblock {On Simplification of Database Integrity Constraints}.
\newblock {\em Fundamenta Informaticae}, 71(4):371--417, 2006.

\bibitem{CM:LAAIC2006}
H.~Christiansen and D.~Martinenghi.
\newblock On using simplification and correction tables for integrity
  maintenance in integrated databases.
\newblock In {\em Proceedings of the Second International Workshop on Logical
  Aspects and Applications of Integrity Constraints (LAAIC06) 8 September 2006,
  Krakow, Poland}, pages 569--576. IEEE Computer Society, 2006.

\bibitem{lncs_2582}
H.~Decker.
\newblock Historical and computational aspects of paraconsistency in view of
  the logic foundation of databases.
\newblock In L.~Bertossi, G.~Katona, K.-D. Schewe, and B.~Thalheim, editors,
  {\em Semantics in Databases}, volume 2582 of {\em LNCS}, pages 63--81.
  Springer, 2001.

\bibitem{Dec02}
H.~Decker.
\newblock Translating advanced integrity checking technology to sql.
\newblock In J.~H. Doorn and L.~C. Rivero, editors, {\em Database integrity:
  challenges and solutions}, pages 203--249. Idea Group Publishing, 2002.

\bibitem{DC94}
H.~Decker and M.~Celma.
\newblock A slick procedure for integrity checking in deductive databases.
\newblock In P.~{Van Hentenryck}, editor, {\em Logic Programming: Proc.\ of the
  11th International Conference on Logic Programming}, pages 456--469. MIT
  Press, Cambridge, MA, 1994.

\bibitem{DM:FlexDBIST2006}
H.~Decker and D.~Martinenghi.
\newblock Avenues to flexible data integrity checking.
\newblock In {\em Proceedings of the International Workshop on Flexible
  Database and Information System Technology (FlexDBIST-06) 6 September 2006,
  Krakow, Poland}, pages 425--429. IEEE Computer Society, 2006.

\bibitem{DM:SEBD2006}
H.~Decker and D.~Martinenghi.
\newblock {Can Integrity Tolerate Inconsistency? (Extended Abstract)}.
\newblock In {\em Proceedings of the Fourteenth Italian Symposium on Advanced
  Database Systems, SEBD 2006, Portonovo, Italy, June 18-21, 2006}, pages
  32--39, 2006.

\bibitem{DM:ADVIS2006}
H.~Decker and D.~Martinenghi.
\newblock Checking violation tolerance of approaches to database integrity.
\newblock In {\em Advances in Information Systems, 4th International
  Conference, ADVIS 2006, Izmir, Turkey, October 18-20, 2006, Proceedings},
  volume 4243 of {\em Lecture Notes in Computer Science}, pages 139--148.
  Springer, 2006.

\bibitem{DM:TDM2006}
H.~Decker and D.~Martinenghi.
\newblock Integrity checking for uncertain data.
\newblock In {\em Proceedings of the Second Twente Data Management Workshop on
  Uncertainty in Databases 6 June 2006, Enschede, The Netherlands}, CTIT
  Workshop Proceedings Series WP06-01, pages 41--48. University of Twente,
  2006.

\bibitem{DM:LPAR2006}
H.~Decker and D.~Martinenghi.
\newblock A relaxed approach to integrity and inconsistency in databases.
\newblock In {\em Logic for Programming, Artificial Intelligence, and
  Reasoning, 13th International Conference, LPAR 2006, Phnom Penh, Cambodia,
  November 13-17, 2006, Proceedings}, volume 4246 of {\em Lecture Notes in
  Computer Science}, pages 287--301. Springer, 2006.

\bibitem{DM:FlexDBIST2007}
H.~Decker and D.~Martinenghi.
\newblock Getting rid of straitjackets for flexible integrity checking.
\newblock In {\em Proceedings of the 2nd International Workshop on Flexible
  Database and Information System Technology (FlexDBIST-07)}, pages 360--364,
  2007.

\bibitem{DM:PPDP2008}
H.~Decker and D.~Martinenghi.
\newblock Classifying integrity checking methods with regard to inconsistency
  tolerance.
\newblock In {\em Proceedings of the 10th International ACM SIGPLAN Conference
  on Principles and Practice of Declarative Programming, July 15-17, 2008,
  Valencia, Spain}, pages 195--204, 2008.

\bibitem{DM:IGP2008}
H.~Decker and D.~Martinenghi.
\newblock Database integrity checking.
\newblock In M.~Khosrow-Pour, editor, {\em Encyclopedia of Information Science
  and Technology (Second Edition)}, volume~II, pages 961--966. Information
  Science Reference, 2008.

\bibitem{DM:IGP2009}
H.~Decker and D.~Martinenghi.
\newblock Inconsistency-tolerant integrity checking.
\newblock In V.~Ferraggine, J.~Doorn, and L.~Rivero, editors, {\em Handbook of
  Research on Innovations in Database Technologies and Applications: Current
  and Future Trends}, volume~II, chapter XXXVIII, pages 348--357. Information
  Science Reference, 2009.

\bibitem{DM:QOIS2009}
H.~Decker and D.~Martinenghi.
\newblock Modeling, measuring and monitoring the quality of information.
\newblock In {\em Proceedings of the 4th International Workshop on Quality of
  Information Systems (QoIS 2009)}, pages 212--221, 2009.

\bibitem{DM:TKDE2011}
H.~Decker and D.~Martinenghi.
\newblock {Inconsistency-tolerant Integrity Checking}.
\newblock {\em IEEE Transactions on Knowledge \& Data Engineering},
  23(2):218--234, 2011.

\bibitem{D76}
E.~W. Dijkstra.
\newblock {\em A Discipline of Programming}.
\newblock Prentice-Hall, 1976.

\bibitem{DBLP:journals/sigmod/DongS00}
G.~Dong and J.~Su.
\newblock {Incremental Maintenance of Recursive Views Using Relational
  Calculus/SQL.}
\newblock {\em SIGMOD Record}, 29(1):44--51, 2000.

\bibitem{Florescu:1998:QCC}
D.~Florescu, A.~Levy, and D.~Suciu.
\newblock Query containment for conjunctive queries with regular expressions.
\newblock In {\em {PODS} '98. Proceedings of the {ACM} {SIGACT--SIGMOD--SIGART}
  Symposium on Principles of Database Systems, June 1--3, 1998, Seattle,
  Washington}, pages 139--148, New York, NY 10036, USA, 1998. ACM Press.

\bibitem{DBLP:conf/sebd/FraternaliMT12}
P.~Fraternali, D.~Martinenghi, and M.~Tagliasacchi.
\newblock Efficient diversification of top-k queries over bounded regions.
\newblock In {\em SEBD}, pages 139--146, 2012.

\bibitem{FMT:SIGMOD2012}
P.~Fraternali, D.~Martinenghi, and M.~Tagliasacchi.
\newblock {Top-k bounded diversification}.
\newblock In {\em Proceedings of the 2012 ACM SIGMOD/PODS Conference -- SIGMOD
  2012, Scottsdale, Arizona, USA, May 20--24, 2012}, pages 421--432, 2012.

\bibitem{GM90}
J.~Grant and J.~Minker.
\newblock Integrity constraints in knowledge based systems.
\newblock In H.~Adeli, editor, {\em Knowledge Engineering Vol II,
  Applications}, pages 1--25. McGraw-Hill, 1990.

\bibitem{310709}
A.~Gupta and I.~S. Mumick, editors.
\newblock {\em Materialized views: techniques, implementations, and
  applications}.
\newblock MIT Press, 1999.

\bibitem{HMN84}
L.~Henschen, W.~McCune, and S.~Naqvi.
\newblock Compiling constraint-checking programs from first-order formulas.
\newblock In H.~Gallaire, J.~Minker, and J.-M. Nicolas, editors, {\em Advances
  In Database Theory, February 1-5, 1988, Los Angeles, California, USA},
  volume~2, pages 145--169. Plenum Press, New York, 1984.

\bibitem{H69}
C.~Hoare.
\newblock An axiomatic basis for computer programming.
\newblock {\em Communications of the ACM}, 12(10):576--580, 1969.

\bibitem{DBLP:conf/sigmod/HsuI85}
A.~Hsu and T.~Imielinski.
\newblock Integrity checking for multiple updates.
\newblock In S.~B. Navathe, editor, {\em Proceedings of the 1985 ACM SIGMOD
  International Conference on Management of Data, Austin, Texas, May 28-31,
  1985}, pages 152--168. ACM Press, 1985.

\bibitem{IMT:SECO2009}
I.~F. Ilyas, D.~Martinenghi, and M.~Tagliasacchi.
\newblock Rank-join algorithms for search computing.
\newblock In S.~Ceri and M.~Brambilla, editors, {\em Search Computing:
  Challenges and Directions}, pages 211--224. 2009.

\bibitem{kowalski_book}
R.~A. Kowalski.
\newblock {\em Logic for Problem Solving}.
\newblock Elsevier, 1979.

\bibitem{KSS87}
R.~A. Kowalski, F.~Sadri, and P.~Soper.
\newblock Integrity checking in deductive databases.
\newblock In P.~M. Stocker, W.~Kent, and P.~Hammersley, editors, {\em VLDB
  1987, Proceedings of 13th International Conference on Very Large Data Bases,
  September 1-4, 1987, Brighton, England}, pages 61--69. Morgan Kaufmann, 1987.

\bibitem{LL96}
S.~Y. Lee and T.~W. Ling.
\newblock Further improvements on integrity constraint checking for
  stratifiable deductive databases.
\newblock In T.~M. Vijayaraman, A.~P. Buchmann, C.~Mohan, and N.~L. Sarda,
  editors, {\em VLDB'96, Proceedings of 22th International Conference on Very
  Large Data Bases, September 3-6, 1996, Mumbai (Bombay), India}, pages
  495--505. Morgan Kaufmann, 1996.

\bibitem{LD98}
M.~Leuschel and D.~de~Schreye.
\newblock Creating specialised integrity checks through partial evaluation of
  meta-interpreters.
\newblock {\em Journal of Logic Programming}, 36(2):149--193, 1998.

\bibitem{LST87}
J.~W. Lloyd, L.~Sonenberg, and R.~W. Topor.
\newblock Integrity constraint checking in stratified databases.
\newblock {\em Journal of Logic Programming}, 4(4):331--343, 1987.

\bibitem{LT85}
J.~W. Lloyd and R.~W. Topor.
\newblock {A basis for deductive database systems II}.
\newblock {\em Journal of Logic Programming}, 30(1):55--67, 1986.

\bibitem{martinenghi2003simplification}
D.~Martinenghi.
\newblock A simplification procedure for integrity constraints.
\newblock \url{http://www.dat.ruc.dk/~dm/spic/index.html}, 2003.

\bibitem{M:ADBIS2004}
D.~Martinenghi.
\newblock Optimal database locks for efficient integrity checking.
\newblock In {\em Eighth East-European Conference on Advances in Databases and
  Information Systems (ADBIS 2004), Budapest, Hungary, 22-25 September 2004,
  local proceedings}, pages 64--77, 2004.

\bibitem{M:FQAS2004}
D.~Martinenghi.
\newblock Simplification of integrity constraints with aggregates and
  arithmetic built-ins.
\newblock In {\em Flexible Query Answering Systems, 6th International
  Conference, FQAS 2004, Lyon, France, June 24-26, 2004, Proceedings}, volume
  3055 of {\em Lecture Notes in Computer Science}, pages 348--361. Springer,
  2004.

\bibitem{M:PHD2005}
D.~Martinenghi.
\newblock {\em Advanced Techniques for Efficient Data Integrity Checking}.
\newblock PhD thesis, Roskilde University, Denmark, in {\em Datalogiske
  Skrifter}, 105, \url{http://www.ruc.dk/dat/forskning/skrifter/DS105.pdf},
  2005.

\bibitem{M:CRYPT2011}
D.~Martinenghi.
\newblock Access pattern.
\newblock In H.~C. van Tilborg and S.~Jajodiathe, editors, {\em Encyclopedia of
  Cryptography and Security (Second Edition)}, pages A17--A20. Springer, 2011.

\bibitem{M:CORR2013b}
D.~Martinenghi.
\newblock On the dependency on the size of the data when chasing under
  conceptual dependencies.
\newblock {\em CoRR}, submit/0863729, 2013.

\bibitem{MC:ADBIS2005}
D.~Martinenghi and H.~Christiansen.
\newblock Efficient integrity checking for databases with recursive views.
\newblock In {\em Advances in Databases and Information Systems, 9th East
  European Conference, ADBIS 2005, Tallinn, Estonia, September 12-15, 2005,
  Proceedings}, volume 3631 of {\em Lecture Notes in Computer Science}, pages
  109–--124. Springer, 2005.

\bibitem{MC:DEXA2005}
D.~Martinenghi and H.~Christiansen.
\newblock Transaction management with integrity checking.
\newblock In {\em 16th International Conference on Database and Expert Systems
  Applications}, volume 3588 of {\em Lecture Notes in Computer Science}, pages
  606–--615. Springer, 2005.

\bibitem{MCD:IGP2006}
D.~Martinenghi, H.~Christiansen, and H.~Decker.
\newblock Integrity checking and maintenance in relational and deductive
  databases - and beyond.
\newblock In Z.~Ma, editor, {\em Intelligent Databases: Technologies and
  Applications}, chapter~X, pages 238--285. Idea Group Publishing, 2006.

\bibitem{MT:PVLDB2010}
D.~Martinenghi and M.~Tagliasacchi.
\newblock {Proximity Rank Join}.
\newblock {\em Proceedings of the VLDB Endowment}, 3(1):352--363, 2010.

\bibitem{MT:DBRANK2010}
D.~Martinenghi and M.~Tagliasacchi.
\newblock Top-k pipe join.
\newblock In {\em Proceedings of the 4th International Workshop on Ranking in
  Databases (DBRank 2010)}, pages 16--19. IEEE Computer Society Press, 2010.

\bibitem{MT:TKDE2011}
D.~Martinenghi and M.~Tagliasacchi.
\newblock {Cost-Aware Rank Join with Random and Sorted Access}.
\newblock {\em IEEE Transactions on Knowledge \& Data Engineering},
  24(12):2143--2155, 2012.

\bibitem{MT:TODS2012}
D.~Martinenghi and M.~Tagliasacchi.
\newblock {Proximity measures for rank join}.
\newblock {\em ACM Transactions on Database Systems}, 37(1), 2012.
\newblock Extended version of~\cite{MT:PVLDB2010}.

\bibitem{MT:ITAIS2009}
D.~Martinenghi and R.~Torlone.
\newblock A logical approach to context-aware databases.
\newblock In {\em Proceedings of the 6th Conference of the Italian Chapter of
  AIS (itAIS 2009)}, 2009.
\newblock Post-proceedings published in~\cite{MT:ITAIS2010}.

\bibitem{MT:TECHREP2009}
D.~Martinenghi and R.~Torlone.
\newblock A model and a language for context-aware databases.
\newblock Technical Report RT-DIA-152-2009, Dipartimento di Informatica e
  Automazione, Universit\`a degli studi Roma Tre, July 2009.

\bibitem{MT:FQAS2009}
D.~Martinenghi and R.~Torlone.
\newblock Querying context-aware databases.
\newblock In {\em Flexible Query Answering Systems, 8th International
  Conference, FQAS 2009, Roskilde, Denmark, October 26-28, 2009. Proceedings},
  pages 76--87, 2009.

\bibitem{MT:ITAIS2010}
D.~Martinenghi and R.~Torlone.
\newblock A logical approach to context-aware databases.
\newblock In A.~D'Atri, M.~D. Marco, A.~Braccini, and F.~Cabiddu, editors, {\em
  Management of the Interconnected World}, pages 211--220. 2010.
\newblock Extended version of~\cite{MT:ITAIS2009}.

\bibitem{MT:ER2010}
D.~Martinenghi and R.~Torlone.
\newblock {Querying Databases with Taxonomies}.
\newblock In {\em Proceedings of Conceptual Modeling - ER 2010, 29th
  International Conference on Conceptual Modeling, Vancouver, BC, Canada,
  November 1-4, 2010}, pages 377--390, 2010.

\bibitem{Nic82}
J.-M. Nicolas.
\newblock Logic for improving integrity checking in relational data bases.
\newblock {\em Acta Informatica}, 18:227--253, 1982.

\bibitem{DBLP:conf/vldb/Olive91}
A.~Oliv{\'e}.
\newblock Integrity constraints checking in deductive databases.
\newblock In G.~M. Lohman, A.~Sernadas, and R.~Camps, editors, {\em 17th
  International Conference on Very Large Data Bases, September 3-6, 1991,
  Barcelona, Catalonia, Spain, Proceedings}, pages 513--523. Morgan Kaufmann,
  1991.

\bibitem{Qia88}
X.~Qian.
\newblock An effective method for integrity constraint simplification.
\newblock In {\em Proceedings of the Fourth International Conference on Data
  Engineering, February 1-5, 1988, Los Angeles, California, USA}, pages
  338--345. IEEE Computer Society, 1988.

\bibitem{KH:VLDBJ94}
K.~Ramamohanarao and J.~Harland.
\newblock An introduction to deductive database languages and systems.
\newblock {\em VLDB Journal}, 3(2):107--122, 1994.

\bibitem{RMT:LID2011}
A.~Rauseo, D.~Martinenghi, and L.~Tanca.
\newblock Context through answer set programming (abstract).
\newblock In {\em Proceedings of the EDBT 2011 workshops, 3rd International
  Workshop on Logic in Databases (LID 2011)}, page~58, 2011.

\bibitem{SK88}
F.~Sadri and R.~Kowalski.
\newblock A theorem-proving approach to database integrity.
\newblock In J.~Minker, editor, {\em Foundations of Deductive Databases and
  Logic Programming}, pages 313--362. Morgan Kaufmann, Los Altos, CA, 1988.

\bibitem{S95}
R.~Selj{\'e}e.
\newblock A new method for integrity constraint checking in deductive database.
\newblock {\em Data \& Knowledge Engineering}, 15(1):63--102, 1995.

\bibitem{SS99}
R.~Selj{\'e}e and H.~C.~M. de~Swart.
\newblock Three types of redundancy in integrity checking: An optimal solution.
\newblock {\em Data \& Knowledge Engineering}, 30(2):135--151, 1999.

\bibitem{SIMT:SIGMOD2011}
M.~A. Soliman, I.~F. Ilyas, D.~Martinenghi, and M.~Tagliasacchi.
\newblock {Ranking with Uncertain Scoring Functions: Semantics and Sensitivity
  Measures}.
\newblock In {\em Proceedings of the 2011 ACM SIGMOD/PODS Conference -- SIGMOD
  2011, Athens, Greece, June 12--16, 2011}, pages 805--816, 2011.

\bibitem{vanMeyden_survey}
R.~van~der Meyden.
\newblock Logical approaches to incomplete information: A survey.
\newblock In {\em Logics for Databases and Information Systems}, pages
  307--356, 1998.

\bibitem{Vardi86}
M.~Y. Vardi.
\newblock On the integrity of databases with incomplete information.
\newblock In {\em PODS}, pages 252--266, 1986.

\end{thebibliography}

\documentclass{article}
\pdfoutput=1
\usepackage{latexsym}
\usepackage{amssymb}
\usepackage{url}
\usepackage{colortbl}

\def\false{\mbox{\textit{false}}}
\def\iff{\mbox{ iff }}
\def\metaequiv{\iff}

\def\pre{\mbox{pre}}
\def\post{\mbox{post}}

\def\true{\mbox{\textit{true}}}

\newtheorem{definition}{Definition}[section]

\newtheorem{example}{Example}[section]

\newtheorem{proposition}[definition]{Proposition}
\newenvironment{proof}{\noindent\textbf{Proof}\itshape}{~\hfill~$\Box$ \newline\newline}
\newtheorem{corollary}[definition]{Corollary}

\title{On the difference between checking\\ integrity constraints before or after updates}

\begin{document}
\author{Davide Martinenghi\\\\
Politecnico di Milano\\
Piazza Leonardo da Vinci 32\\
20133 Milano, Italy\\
\url{davide.martinenghi@polimi.it}}
\date{}

\maketitle

\begin{abstract}
Integrity checking is a crucial issue, as databases change their instance all the time and therefore need to be checked continuously and rapidly.
Decades of research have produced a plethora of methods for checking integrity constraints of a database in an incremental manner.
However, not much has been said about \emph{when} to check integrity.
In this paper, we study the differences and similarities between checking integrity before an update (a.k.a. pre-test) or after (a.k.a. post-test) in order to assess the respective convenience and properties.
\end{abstract}

\section{Introduction}\label{sec:intro}

Integrity checking has been 
a perennial topic in almost all database conferences, journals and
research labs. Very large quantities of research activities and
publications testify to the importance of the issue. The motivation
which has stimulated these investigations is that integrity checking
is practically unfeasible for significant amounts of stored data
without a dedicated approach to optimize the process.
Progress has been made with extensions of basic approaches to
deductive, object-relational, XML-based, distributed and other kinds
of advanced database technology, as surveyed in \cite{MCD:IGP2006}. However,
what has not changed much are the fundamental ideas that are already
present in the seminal paper \cite{Nic82}.

The basic idea is that, in most cases, a simplified form of the set
of integrity constraints imposed on the database can be obtained
from a given update (or just an update schema) and the current state
of the database (or just the database schema). Thus, integrity,
which is supposed to be an invariant of all possible database
states,
is checked upon each update request, which in turn becomes effective
if the check yields that integrity is not violated. Here,
``simplified'' essentially means more efficiently evaluated at update
time.
Of course, efficiency is not unequivocally measurable. However, the
number of stored facts to be retrieved for constraint evaluation is
a good rule of thumb. Another such rule is the number of literals in
the simplified form. Also, the minimization (reduction or avoidance)
of repair costs
after having detected integrity violation is a factor to be
considered when assessing the efficiency of integrity checking.
To establish the new state is usually considered less costly than to
undo it, but for concurrent transactions, more so in distributed and
even more so in replicated databases, establishing the new state is
a sizable and non-negligible cost factor (think of the concurrency
control, management and communication rounds needed for distribution
and different replication strategies), while roll-backs are less of
a problem since they can be taken care of and optimized by the DBMS,
making use of its transaction logs.

Integrity checking methods may differ in some of the assumptions.
For instance, in \cite{Nic82} and in most of the publications on the same
subject that came after it, a categorical premise for the
correctness of the simplification approach has been that the
constraints to be checked 
for a given update $U$ are supposed to be satisfied in the ``old"
state, i.e., the database state given when $U$ is requested.
This assumption has been relaxed in~\cite{MT:TKDE2011}, thereby introducing the notion of inconsistency-tolerant integrity checking.
Other different assumptions may regard the class of integrity constraints and updates that are supported by the methods, as surveyed in~\cite{MCD:IGP2006}.

Unlike previous surveys, in this paper we study the main differences between approaches that require to check integrity after all updates have been applied and those that check integrity before the updates.
We shall see that these checks are generally non-interchangeable, unless the updates are of a specific kind.

\section{Preliminaries}
\label{sec:integrity-constraints-simplified}

In this section, we recapitulate, classify and discuss the main
characteristics of the simplification approach to integrity
checking. We fix basic
definitions and terminology, and
thus the framework into which our formalizations in the remainder
are cast. In Section \ref{subsec:simplifications} we introduce
abstract notions of various classes of simplifications and discuss
them in Section \ref{subsec:pre-tests-vs-post-tests}.

Throughout, we refer to the relational framework of deductive
databases,
i.e., relational databases with possibly recursive view definitions
described in clause form \cite{AHV95,KH:VLDBJ94}.
Thus, a {\em database} consists of a set of \emph{facts}
and a set of \emph{rules}, i.e.,
\emph{tuples} and \emph{views}, 
in the terminology of the relational model.

An {\em integrity constraint} expresses a semantic invariant, i.e.,
a condition supposed to hold in each state of the database. In
general, it can be expressed by any closed first-order logic formula
in the language of the database on which it is imposed.
Two kinds of normal form representations of integrity constraints
which both incur no loss of generality are prominent in the
literature: prenex normal form and denial form. The former, as
defined and used, e.g., in \cite{Nic82, LST87}, has all quantifiers
moved leftmost and all negation symbols moved innermost, by
equivalent rewritings of the original formula. The latter, as
defined and used, e.g., in \cite{SK88,KSS87}, has the form of
datalog clauses with empty head, expressing that, if their condition
is satisfied, then integrity is violated, and may need dedicated
view definitions to define these conditions by recurring on database
facts.
An 
\emph{integrity theory} 
is a finite set of integrity constraints, 
to be thought of as being imposed on some database.

For simplicity, we limit ourselves to databases that have a
unique standard model and no ``unknown" facts (i.e., each fact is
either true or false in the model), e.g., stratified databases,
and assume that database semantics is defined 
by this model.

For a closed formula $W$, we write $D\models W$ (resp.,
$D\not\models W$) to indicate that $W$ evaluates to $\true$ (resp.,
$\false$) in $D$'s standard model.
For a set of formulas $\Gamma$, we write $D\models \Gamma$ (resp.,
$D\not\models \Gamma$) to indicate that for every (resp., some)
formula $W\in\Gamma$ we have $D\models W$ (resp., $D\not\models W$).
If $W$ is an integrity constraint and $\Gamma$ an integrity theory,
it is usual to also say that $D$ {\em satisfies} (resp., {\em
violates}) $W$ and $\Gamma$, respectively. An equivalent terminology
is defined as follows.
\begin{definition}[Consistency]
A database $D$ is {\em consistent} with an integrity theory $\Gamma$
iff $D\models\Gamma$.
\end{definition}
Informally, we have already spoken of database states. More
formally, database states are determined by atomically executed
updates. An \emph{update} $U$ is a mapping $U:{\cal D} \mapsto {\cal
D}$, where ${\cal D}$ is the space of databases
as determined by a
fixed, sufficiently rich language which need 
not be extended by any update. For simplicity, we only consider
updates that may involve facts and rules in this paper, but no
integrity constraints, which are thought of as immutable.
For convenience, for any database $D$,
let $D^U$ denote the new database 
obtained by 
applying update $U$ on $D$. 

\section{Kinds of simplifications}
\label{subsec:simplifications}

Traditionally,
the integrity checking problem asks, given a set of integrity
constraints $\Gamma$, a database $D$ consistent with $\Gamma$, and
an update $U$,
whether $D^U\models\Gamma$ holds, i.e., whether the new database is
consistent with (i.e., satisfies) the integrity constraints.
However,
evaluating $\Gamma$ in $D^U$ may be prohibitively expensive. 
So, a reformulation of the problem is called for, trying to take
advantage of the incrementality of updates. Traditionally, all such
reformulations have been made
under the assumption that the old state 
is consistent.

We will discuss two kinds of 
such reformulations 
that have commonly been 
dealt with in the literature.
Both determine an alternative integrity theory $\Upsilon$ (which by
itself is later called a simplification), the evaluation of which is
supposed to be simpler than to evaluate $\Gamma$, while yielding
equivalent results.
The first kind 
of such $\Upsilon$
is determined to be evaluated in the new 
state. 
\begin{definition}[Post-test]\label{def:post-test}
Let $\Gamma$ be an integrity theory and $U$ an update. An integrity
theory $\Upsilon$ is a \emph{post-test} of $\Gamma$ for $U$ whenever
$D^U\models\Gamma\mbox{ iff } D^U\models \Upsilon$ for 
every
database $D$ consistent with $\Gamma$.
\end{definition}
Clearly, $\Gamma$ itself is a post-test of $\Gamma$
for 
any update.
As indicated, one is interested in producing a post-test that is
actually ``simpler'' to evaluate than the original integrity
constraints. This
is traditionally 
achieved by exploiting
the fact 
that the old 
state $D$ is consistent with $\Gamma$, thus 
avoiding redundant checks
of cases that are already known to satisfy integrity.
Note that the process of integrity checking involving a post-test
consists in: executing the update, checking the post-test and, if it
fails to hold, correcting the database by
performing a repair action, i.e., a rollback 
and optionally a modification of the update
which won't violate integrity. 
Well-known post-test-based approaches are described in
\cite{Nic82,LT85,GM90,DBLP:conf/vldb/Olive91,LD98}.

The second kind of approach to deal with integrity checking
incrementally is to determine an integrity theory $\Sigma$
to be evaluated in the old state, i.e., to predict without actually
executing the update whether the new, updated state will be
consistent with the integrity constraints.
\begin{definition}[Pre-test]\label{def:pre-test}
Let $\Gamma$ be an integrity theory and $U$ an update. An integrity
theory $\Sigma$ is a \emph{pre-test} of $\Gamma$ for $U$ whenever
$D^U\models\Gamma\mbox{ iff } D\models \Sigma$ for 
every
database $D$ consistent with $\Gamma$.
\end{definition}
%
Here, not only the consistency of the old state $D$ with $\Gamma$ is
exploited, but also that inconsistent states can be prevented
without executing the update, and, thus, without ever
having to undo a violated new state. 
The integrity checking process involving a pre-test is therefore:
check whether the pre-test holds and, if it does, execute the
update. Examples of pre-test-based approaches are
\cite{Qia88,CM:FI2006}.
Note that, depending on the requirements of availability and
consistency of a given application, integrity checking with
pre-tests is possibly better suited for concurrent transaction
processing, particularly in distributed databases.

In the remainder, we refer to both post- 
and pre-tests 
as \emph{simplifications} of the original integrity theory.
It is tacitly assumed that
given simplifications
are, at least in significant classes of cases, indeed simpler to
evaluate than the original constraints. A characterization of
simplicity beyond what is mentioned in the introduction, e.g., by
formal cost models,
is out of the scope of this paper;
cf.~\cite{M:PHD2005} 
for a discussion.

Now, whatever the definition of simplicity, 
we cannot directly compare the evaluation cost
of a post-test 
with that of a pre-test 
since they refer to two different (viz. old and new) states.
Therefore,
it is desirable to have kindred 
reference pre- and post-tests for benchmarking given
simplifications.
Plain 
tests, as defined below, i.e., simplifications that do not exploit
the fact that the old state satisfies integrity,
may serve as such reference tests.
\begin{definition}[Plain test]\label{def:plain-test}
Let $\Gamma$ be an integrity theory and $U$ an update.

a) An integrity theory $\Sigma_0$ is a \emph{plain pre-test} of
$\Gamma$ for
$U$, denoted by 
$\pre_0^U(\Gamma)$, if the following holds: $D\models\Sigma_0 \mbox{
iff } D^U\models\Gamma$ for every database $D$.

b) An integrity theory $\Upsilon_0$ is a \emph{plain post-test} of
$\Gamma$ for $U$, denoted by $\post_0^U(\Gamma)$, if the following
holds: $D^U\models\Upsilon_0 \mbox{ iff } D^U\models\Gamma$  for
every database $D$.
\end{definition}
Clearly, $\Gamma$ is a plain 
post-test
 of itself
for 
any update.
Note that each 
plain test is also a simplification.
For any pre-test (resp., post-test),
it is therefore desirable that it be at least as simple
to evaluate 
as the corresponding plain test. 

It is straightforward to see that, for fixed $\Gamma$ and $U$, all
plain 
pre-tests of $\Gamma$ for 
$U$ are logically equivalent.
\begin{proposition}\label{pro:all-plain-pre-tests-equivalent}
Let $\Gamma$ be an integrity theory and $U$ an update. Then, for any
two 
plain
pre-tests $\Sigma'$ and $\Sigma''$ of $\Gamma$ for
$U$, we have $\Sigma'\equiv\Sigma''$.
\end{proposition}
\begin{proof}
By applying definition \ref{def:plain-test} to $\Sigma'$ and
$\Sigma''$, one gets by transitivity that $D\models\Sigma'\iff
D\models\Sigma''$ for 
every
$D$, i.e., $\Sigma'\equiv\Sigma''$.
\end{proof}
Conversely, not all plain post-tests are logically equivalent.
As a counterexample,
take, e.g., $\Gamma=\{\leftarrow p(a)\}$ and let $U$ be the
insertion of $p(a)$. Then $\Upsilon=\false$ is a post-test of
$\Gamma$ for 
$U$ but 
$\Upsilon\not\equiv\Gamma$.

We conclude this section with an example of simplification of
integrity constraints.
\begin{example}\label{ex:conflict-of-interests}
Consider a database with the relations $rev(S,R)$ (submission $S$
assigned to reviewer $R$), $sub(S,A)$ (submission $S$ authored by
$A$) and $pub(P,A)$ (publication $P$ authored by $A$). Assume a
policy establishing that no one can review a paper of his/her
(former) coauthors. This is expressed by:
$$\begin{array}{rl}
\!\!\Gamma\!\!=\!\!\{&\!\!\!\!\leftarrow \!\!rev(S,R)\land sub(S,R),\\
\!\!& \!\!\!\!\leftarrow \!\!rev(S,R)\land sub(S,A)\land
pub(P,R)\land pub(P,A)\}
\end{array}$$
Let $U$ be the an update that inserts the facts $sub(c,a)$ and
$rev(c,b)$ into the database, where $a$, $b$, $c$ are some
constants.
A simplification of $\Gamma$ for $U$ (equivalent to what Nicolas'
method would output) is as follows:
$$\begin{array}{rl}
\Sigma=\{&\leftarrow sub(c,b),\\
& \leftarrow rev(c,a),\\
& \leftarrow pub(P,b)\land pub(P,a),\\
& \leftarrow sub(c,A)\land pub(P,b)\land pub(P,A),\\
& \leftarrow rev(c,R)\land pub(P,R)\land pub(P,a)\;\;\}
\end{array}$$
The simplified conditions given by $\Sigma$ can be read as follows:
\begin{itemize}
\item $b$ did not submit $c$
\item $a$ does not review $c$
\item $b$ is not coauthor of $a$
\item $b$ is not coauthor of an author of $c$
\item $c$ is not reviewed by a coauthor of $a$
\end{itemize}
These checks are much easier to execute than $\Gamma$, as they
greatly reduce the space of tuples to be considered by virtue of the
instantiation of variables with constants.
\end{example}

\section{Relationship between pre- and post-tests}
\label{subsec:pre-tests-vs-post-tests}
In this section we compare pre-tests and post-tests and discuss
their interchangeability.
Note that 
we do this without referring to any specific simplification method
or update language.

First, we show that, in general, a pre-test cannot be used as a
post-test, nor vice versa.

\begin{example}\label{ex:pre-and-post-test-differ-in-general}
Consider the integrity theory $\Gamma=\{\leftarrow p(a)\land q(b)\}$
and an update $U$ that exchanges the contents of $p$ and $q$. Then
$\Sigma=\{\leftarrow q(a)\land p(b)\}$ is a pre-test but clearly not
a correct post-test. Consider, e.g., a database
$D=\{p(a),p(b),q(a)\}$; we have $D\models\Gamma,
D^U\not\models\Gamma, D^U\models\Sigma$, i.e., it does not hold that
$D^U\models\Sigma \metaequiv D^U\models\Gamma$, although $D$ is
consistent with $\Gamma$. Similarly, $\Upsilon=\{\leftarrow
p(a)\land q(b)\}$ is a post-test, but not a pre-test, of $\Gamma$
for $U$.
\end{example}
This result is not surprising, since we allow for updates
representing any kind of transformation of the database, such as
swapping the contents of two relations.
We now introduce a class of updates that excludes an update such as
$U$ of example \ref{ex:pre-and-post-test-differ-in-general}.
\begin{definition}\label{def:idempotent-update}
An update $U$ is \emph{idempotent}\index{update!idempotent update}
if $D^U$ = $(D^U)^U$ for any database $D$.
\end{definition}
Idempotent updates capture additions, deletions and changes of
specific tuples, which are certainly among the most frequent kinds
of updates.
For idempotent updates, we can prove that a plain pre-test is also
always a valid plain post-test.
\begin{proposition}\label{pro:plain-pre-test=post-test-idempotent}
Let $\Gamma$ be an integrity theory and $U$ an idempotent update.
Then $D^U\models\pre_0^U(\Gamma)\metaequiv D^U\models\Gamma$ for any
database $D$, i.e., $pre_0^U(\Gamma)$ is a plain post-test of
$\Gamma$ for $U$.
\end{proposition}
\begin{proof}
Since $U$ is idempotent, i.e., $D^U = (D^U)^U$ for any $D$, we have
\begin{itemize}
\item[(1)] $D^U\models\Gamma\metaequiv (D^U)^U\models\Gamma$ for any
$D$.
\end{itemize}
Since $\pre_0^U(\Gamma)$ is a plain pre-test of $\Gamma$ wrt. $U$,
we have
\begin{itemize}
\item[(2)] $D^U\models\pre_0^U(\Gamma)\metaequiv
(D^U)^U\models\Gamma$ for any $D^U$ (and thus for any $D$).
\end{itemize}
\noindent By transitivity between (1) and (2) we obtain the thesis.
\end{proof}
More surprisingly, however, the converse does not hold, i.e., there
are plain post-tests that are not plain pre-tests. In general, even
for idempotent updates, there are pre-tests that are not post-tests
and post-tests that are not pre-tests.
\begin{proposition}\label{pro:pre-differ-post-for-idempotent}
Let $\Gamma$ be an integrity theory and $U$ an idempotent update.
Then
\begin{enumerate}
\item [(1)] there is a pre-test $\Sigma$ of $\Gamma$ for $U$ such
that it does not hold that $D^U\models\Sigma\metaequiv
D^U\models\Gamma$ for any database $D$ consistent with $\Gamma$,
i.e., $\Sigma$ is not a post-test of $\Gamma$ for $U$.
\item [(2)] there is a plain post-test $\Upsilon$ of $\Gamma$ for
$U$ such that it does not hold that $D\models\Upsilon\metaequiv
D^U\models\Gamma$ for any database $D$ consistent with $\Gamma$,
i.e., $\Upsilon$ is not a pre-test of $\Gamma$ for $U$.
\end{enumerate}
\end{proposition}
\begin{proof}

\noindent (1) Let $D$ be a state consistent with $\Gamma$. We have:
$$\begin{array}{rll}
D\models\Sigma&\metaequiv D^U\models \Gamma \mbox{ ($\Sigma$ pre-test)}\\
&\metaequiv (D^U)^U\models \Gamma \mbox{ ($U$ idempotent)
}\\
&\metaequiv D^U\models \Sigma, \mbox{ if $D^U\models\Gamma$
($\Sigma$ pre-test) }
\end{array}$$
So, the only possibility is a situation where $D\not\models\Sigma$,
$D^U\models\Sigma$, and $D^U\not\models\Gamma$, which happens, e.g.,
with $\Gamma = \{\leftarrow p(a)\}$, $\Sigma = \{\leftarrow
\sim$$p(a)\}$, $D=\emptyset$ and $U$ an update such that
$D^U=\{p(a)\}$.
To conclude the proof, we show that the chosen $\Sigma$ is a
pre-test. Indeed, for any $D$ such that $D\models\Gamma$, then
$D\not\models\Sigma$, and $D^U\not\models\Gamma$.

\noindent(2) The same $D$, $U$ and $\Gamma$ as in the previous point
can also be used for the second case by considering the plain
post-test $\Upsilon=\Gamma$. We have $D\models\Gamma$,
$D^U\not\models\Gamma$, $D\models\Upsilon$ and
$D^U\not\models\Upsilon$.
\end{proof}
Another interesting aspect regarding pre-tests and post-tests is
whether their evaluation is at all affected by the update.
Proposition \ref{pro:plain-pre-test=post-test-idempotent}
immediately implies that the evaluation of a plain pre-test is not
affected by the update, as stated in the following corollary.
\begin{corollary}\label{cor:plain-pre-test-resource-set-update-disjoint}
Let $\Gamma$ be an integrity theory and $U$ an idempotent update.
Then $D\models\pre_0^U(\Gamma)\metaequiv D^U\models\pre_0^U(\Gamma)$
for any $D$.
\end{corollary}
\begin{proof}
By definition of $\pre_0^U(\Gamma)$, we have
\begin{itemize}
\item[] $D\models\pre_0^U(\Gamma)\metaequiv D^U\models\Gamma$ for
any $D$,
\end{itemize}
and, by transitivity with the claim of proposition
\ref{pro:plain-pre-test=post-test-idempotent}, we have the thesis.
\end{proof}
However, this does not hold in general for pre-tests or (plain)
post-tests, as demonstrated in the example of the proof of
proposition \ref{pro:pre-differ-post-for-idempotent}.

The following table summarizes the results presented in this
section. We indicate with Pre$^U(\Gamma)$ (resp., Post$^U(\Gamma)$)
the set of all pre-tests (resp., post-tests) of integrity theory
$\Gamma$ for $U$, and use a $0$ subscript to indicate the set of all
plain pre-tests (resp., post-tests) of $\Gamma$ for $U$.
\vspace{0.2cm}

\setlength{\extrarowheight}{0.05cm}
\noindent\begin{center}\begin{tabular}{|l|c|c|} \hline
for any $\Gamma$ & any $U$ & $U$ idempotent\\
\hline
Pre$^U(\Gamma)\subseteq$Post$^U(\Gamma)$? & \textbf{no} & \textbf{no}\\
\hline
Post$^U(\Gamma)\subseteq$Pre$^U(\Gamma)$? & \textbf{no} & \textbf{no}\\
\hline
Pre$_0^U(\Gamma)\subseteq$Pre$^U(\Gamma)$? & \textbf{yes} & \textbf{yes}\\
\hline
Post$_0^U(\Gamma)\subseteq$Post$^U(\Gamma)$? & \textbf{yes} & \textbf{yes}\\
\hline
Pre$_0^U(\Gamma)\subseteq$Post$^U(\Gamma)$? & \textbf{no} & \textbf{yes}\\
\hline
\end{tabular}\end{center}
\setlength{\extrarowheight}{0cm}

\section{Related work}\label{sec:related}
%
Simplification of integrity constraints has been recognized by a
large body of research as a powerful technique for optimization of
integrity checking.
Several approaches to simplification 
require the update transaction to be performed
{\em before} the resulting state is 
checked for consistency with a
post-test~\cite{Nic82,LST87,SK88,GM90,DC94,LD98}.
%
%
As opposed to that,
to pre-test the feasibility of an update with respect to an
integrity theory allows for avoiding both the execution of the
update and, particularly, the restoration of the database state
before the update, which may be very costly.
%
Pre-test-based methods are, e.g.,
\cite{HMN84,DBLP:conf/sigmod/HsuI85,Qia88,LL96,LD98,CM:FI2006,DM:IGP2008,MCD:IGP2006,CM:LPAR2005,MC:ADBIS2005,MC:DEXA2005,M:ADBIS2004,M:FQAS2004,CM:LOPSTR2003,M:PHD2005,martinenghi2003simplification,CM:FoIKS2004,CM:LAAIC2006},
including a few industrial attempts, e.g.,
\cite{Carrico95,Benedikt02}.
Other methods provide simplifications that may require the
availability of both the old and the new state, assuming that the
database keeps track of the old state before committing an update,
\cite{S95,SS99}.

In \cite{GM90}, an adaptation of subsumption checking (called
\emph{partial subsumption}) is used to generate simplification as
the ``difference'' (called \emph{residual}) between an integrity
constraint and a clause representing an update.

Qian's method~\cite{Qia88} generates pre-tests for integrity
checking based on the observation that a simplified integrity
constraint can be regarded as a weakest precondition for having a
consistent updated state, in the same sense as in Hoare's
logic~\cite{H69,D76} for imperative languages, and by assuming
consistency of the database before the update.

Simplification of integrity constraints for update patterns
resembles the notion of program \emph{specialization} used in
partial evaluation, which is the process of creating a specialized
version of a given program (in this case, a general integrity
checker) with respect to known input data (here, the update), as
investigated in \cite{LD98}.

More generally, integrity checking can be seen as a special case of
materialized view maintenance: integrity constraints are defined as
views that must always remain empty for the database to be
consistent~\cite{310709,DBLP:journals/sigmod/DongS00}.

Simplification can also be obtained by resorting to decision
procedures for query
containment~\cite{Florescu:1998:QCC,Calvanese:1998:DQC,CDV03,DBLP:conf/pods/Bonatti04},
as shown in \cite{CM:FI2006}.

We intentionally did not do so before, but at this point is seems
worth mentioning that several simplification methods accept
instantiable or parameterizable {\em patterns} of
updates instead of specific updates, e.g. 
~\cite{HMN84,GM90,CM:FI2006}. Thus, given such a pattern at schema
specification time, it is possible to compile a simplification of
the integrity theory for all updates matching that pattern. For
instance, if constants $a$, $b$ and $c$ in the update of Example
\ref{ex:conflict-of-interests} were specified as simple placeholders
for constants (called \emph{dummy constants} in \cite{HMN84} and
\emph{parameters} in \cite{CM:FI2006}), and thus not assumed to be
necessarily different, a pre-test-based simplification would also
include an integrity constraint that checks that $a\neq b$.

Logic programming-based approaches such as \cite{SK88,KSS87} do not
take into account irrelevant clauses for refuting denial
constraints, even if they would take part in an unnoticed case of
inconsistency that has not been caused by the checked update but by
some earlier event. Moreover, such approaches do not exhibit any
explosive behavior as predicted classical logic in the presence of
inconsistency. In other words, query evaluation procedures based on
SL-resolution can fairly well be called inconsistency-tolerant or
``paraconsistent" in a procedural sense, as done, e.g., in
\cite{kowalski_book,lncs_2582}.
The declarative inconsistency
tolerance of simplifications for improving integrity checking has been studied in several works~\cite{DM:TKDE2011,DM:PPDP2008,DM:LPAR2006,DM:ADVIS2006,DM:SEBD2006,DM:QOIS2009,DM:FlexDBIST2007,DM:TDM2006,DM:FlexDBIST2006,DM:IGP2009}.

The related problem of restoring integrity of a database once
inconsistencies are discovered is tackled by calculating a
\emph{repair}, i.e., a consistent database that is as close as
possible to the original, inconsistent database. Since the seminal
contribution \cite{ABC00}, many authors have studied the problem of
providing consistent answers to queries posed to an inconsistent
database. These techniques certainly add to inconsistency tolerance
in databases, but cannot be directly used to detect inconsistencies
for integrity checking purposes (i.e., by posing integrity
constraints as queries).
Along the same lines, active rules have been considered as a means
to restore a consistent database
\cite{DBLP:conf/vldb/CeriW90,DBLP:journals/tods/CeriFPT94,Dec02}.

Last, we mention work on incomplete databases which also considers
integrity constraints that are not satisfied in a given database
state as something to be dealt with constructively, instead of
banning it from consideration, as most integrity checking methods do
\cite{Vardi86, vanMeyden_survey,CM:TPLP2010,M:CORR2013b}. However, that work is not
interested in integrity checking simplifications that could be used
in such databases. Rather, it is dealing with the issue of
satisfiability and its computational complexity, as related to an
open world assumption by which the space of possible ``closed
worlds" (i.e. database states without missing information) that
would satisfy integrity are studied. The theme of this paper is to
simplify the checking of integrity satisfaction in the presence of
inconsistency, not to ask for the satisfiability of integrity
constraints in the absence of complete information. (Basic
similarities and differences of satisfaction and satisfiability of
integrity are addressed in \cite{bryDeckerManthey88}.)

Relevant new directions of research regard all those areas where integrity constraints are used to characterize useful scenarios in which query answering plays an important role.
Among these, we mention
\emph{i)} \emph{access patterns}, which are constraints indicating which attributes of a relation schema are used as input and which ones are used as output~\cite{CM:APWEB2010,CM:EDBT2010,M:CRYPT2011,CM:ER2008,CM:ICDE2008,CCM:EROW2007,CMC:SEBD2007,CCM:JUCS2009},
\emph{ii)} top-$K$ queries, where the constraints specify a limit on the number of results that the query should return, including constraints on proximity or diversity~\cite{MT:TODS2012,MT:PVLDB2010,CCFMT:TODS2013,FMT:SIGMOD2012,DBLP:conf/sebd/FraternaliMT12,MT:TKDE2011,CMT:TECHREP2009,MT:DBRANK2010,SIMT:SIGMOD2011,IMT:SECO2009},
\emph{iii)} \emph{taxonomies} and context information, which may be used to pose constraint on the granularity of the data and to reason about it~\cite{MT:ER2010,MT:FQAS2009,MT:ITAIS2010,MT:ITAIS2009,MT:TECHREP2009,RMT:LID2011,CM:AAI2000}.

\section{Conclusion}\label{sec:conclusion}

We have discussed and compared the two main abstract families of methods that can be used to incrementally check integrity constraints: pre-tests and post-tests.
These are simplifications to be checked before or, respectively, after the update is executed (while update commitment is supposed to occur only after a successful check).
In order to not only talk about some selected, specific methods, albeit well-known ones, we have characterized declarative and procedural aspects of simplification-based integrity checking in a manner which is largely independent of concrete methods.
Unsurprisingly, pre-tests and post-tests are not interchangeable, not only in terms of the convenience of executing the ones or the others in practical situations (pre-tests may actually be preferred in case of updates to be rejected), but also of their semantic properties.
Somewhat surprisingly, however, their applicability is mostly asymmetric, even for the simple case of idempotent updates.

\bibliographystyle{abbrv}
{\small \bibliography{DM_VLDB06}}
\end{document}